# Elliptic flow of deuterons in ultrarelativistic heavy-ion collisions

### v2

@@ I Introduction and motivation
Other possibilities to discriminate between the two production models include examining the production of very exotic spatially extended and high spin isotope ${}^{4}$i Bazak:2020wjn , and/or performing femtoscopic studies involving correlations of deuterons to protons, another hadrons Mrowczynski:2019yrr , or among themselves Mrowczynski:2021bzy .
@@ III.1 Freeze-out hypersurface
In this paper, we shall denote the rapidity with capital letter in order to distinguish it from the spatial coordinate. As spatial coordinates we use the radial coordinate $r$ and the azimuthal angle $\Theta$, as well as the space-time rapidity
@@
The expansion velocity is parametrised as
@@ III.4 Model calibration
[…]

### v3

@@ I Introduction and motivation
Other possibilities to discriminate between the two production models include examining the production of very exotic spatially extended and high spin isotope 4Li Bazak:2020wjn , and/or performing femtoscopic studies involving correlations of deuterons to protons, another hadrons Mrowczynski:2019yrr , or among themselves Mrowczynski:2021bzy .
@@ III.1 Freeze-out hypersurface
In this paper, we shall denote the rapidity with capital letter in order to distinguish it from the spatial coordinate.
In the position space, we shall use the polar coordinates, $r$ and the azimuthal angle $\Theta$, in the transverse plane. Hyperbolic coordinates map the remaining two directions: the space-time rapidity
@@
Note that Eqs. (6) and (9) just show how the positions in space and the freeze-out hypersurface are parametrised, respectively, with the help of coordinates. They do not yet desribe the actual shape of the hot fireball.
@@
Technically, $R(\Theta)$ sets the maximum value that the coordinate $r$ can attain. The azimuthal spatial anisotropy of the fireball is set by the $\Theta$-dependence of this maximum. The expansion velocity is parametrised as
@@ III.4 Model calibration
At this point we observed that we are unable to reproduce $v_{2}(p_{t})$ reasonably when we stick to the assumption that the freeze-out hypersurface is set by constant $\tau$ and does not depend on $r$. The reason is that $v_{2}(p_{t})$ requires rather strong elliptic flow. Since $\rho_{0}$ parametrizes the transverse rapidity, the parts of the fireball at the edges move with velocities close to that of the light. To see a sizeable effect on the elliptic flow we need a variation in the transverse expansion velocity. Close to the edge, the variation is rather limited since the expansion velocity is close to $c$ and must remain below $c$, The effect of expansion velocity variation is enhanced if the freeze-out hypersurface bends downwards in the $\tau$-$r$ plane, i.e., regions which are farther from the longitudinal symmetry axis freeze-out earlier [^1] The scenario is sometimes referred to as “burning log”.. Technically, this is expressed by the value $s_{2}=-0.02$ fm-2. The obtained anisotropy parameters for centrality classes in which we calculated $v_{2}(p_{t})$ are listed in Table 1.